OpenStudy (anonymous):

How many different 9-member boards of directors can be formed from a group of 14 candidates? I tried 14)13)12)11)10)9)8/14 but I'm not getting the right answer.

OpenStudy (kropot72):

14C9 \[14C9=\frac{14!}{9!5!}=\frac{14\times 13\times 12\times 11\times 10}{5\times 4\times 3\times 2}=you\ can\ calculate\]

OpenStudy (anonymous):

Okay...so 'splain this to me if you would be so kind.

OpenStudy (kropot72):

What don't you understand?

OpenStudy (anonymous):

Why this is so different from what my book said :). Okay, for starters, why are there only 5 terms on top and 4 on the bottom? Why five and four?

OpenStudy (kropot72):

Do you understand what a factorial is? The formula for finding the number of combinations of n different things taken r at a time is \[nCr=\frac{n!}{r!(n-r)!}\]

OpenStudy (anonymous):

No, never got that handy dandy formula. But okay...so you used 5 terms on the top and bottom (the five came from 14-9) but there's only 4 terms on the bottom because you just dropped the times 1 which would have no effect...am i right?

OpenStudy (kropot72):

\[\frac{14!}{9!}=14\times 13\times 12\times 11\times 10\] and \[5!=5\times 4\times 3\times 2\] As you realise the times 1 is unnecessary.

OpenStudy (anonymous):

Okay I think I get how to use it...I dont completely get the why and how but thats just a luxury I guess.

OpenStudy (kropot72):

There is a good explanation of combinations and permutations and the necessary formulas here: http://www.mathsisfun.com/combinatorics/combinations-permutations.html

OpenStudy (anonymous):

Okay i'll check it out. thanks

OpenStudy (kropot72):

You're welcome :)